30. Data sets A and B are dependent . Find the critical value, to, to test the claim that u(d)-0. Use a-0.05. A 40 38 57 53 4 B 38 34 35 45 32 31. Data sets A and B are dependent. Test the claim that the paired sample data is from a population with a mean difference of 0. Use a = 0.05. A 37 35 54 50 38 B 35 31 32 42 29
Answer #1

As per the HOMEWORKLIB POLICY, we are advised to do one question at a time so i am attempting 30th.

30. This would be a paired sample t - test since the samples are dependent. Morever, this would be a two tailed test.

Degrees of freedom = n - 1 = 5 - 1 = 4

For 4 degrees of freedom and 0.05 level of significance, two tailed critical value will be:

t0 = 2.776